# Understanding Memorization in Generative Models via Sharpness in Probability Landscapes

### v2

@@ 1 Introduction
Despite their successes, diffusion models suffer from memorization, where they replicate training samples instead of generating novel outputs (Carlini et al., 2023; Somepalli et al., 2023b). This issue is especially concerning when models are trained on sensitive data, leading to privacy risks (Orrick, 2023; Joseph Saveri, 2023). Addressing memorization is critical for ensuring the responsible deployment of generative models in real-world applications.
@@
Beyond detection, we introduce an inference-time mitigation strategy that reduces memorization by selecting initial diffusion noise from regions of lower sharpness. Our method, Sharpness-Aware Initialization for Latent Diffusion (SAIL), utilizes our sharpness metric to identify initializations that avoid trajectories leading to memorization. By simply adjusting the initial noise, SAIL steers the diffusion process toward smoother probability regions, mitigating memorization without requiring retraining. Unlike prompt modification, which can degrade generation quality, SAIL optimizes noise selection, offering a new avenue for reducing memorization without altering conditioning inputs.
@@
  • We provide a theoretical justification for Wen’s metric through sharpness analysis.
  • We introduce a new sharpness measure, enabling early-stage memorization detection in the diffusion process.
@@ 2 Related works
The memorization behavior of diffusion models (DMs) has been extensively studied (Somepalli et al., 2023b; Carlini et al., 2023; Wen et al., 2024), with prior work examining contributing factors such as prompt conditioning (Somepalli et al., 2023b), data duplication (Carlini et al., 2023; Somepalli et al., 2023a), and dataset size or complexity (Gu et al., 2023). Some studies have approached this issue from a geometric standpoint, drawing on the manifold learning conjecture (Fefferman et al., 2016; Pope et al., 2021), where exact memorization is associated with data points lying on a zero-dimensional manifold (Ross et al., 2024; Ventura et al., 2024; Pidstrigach, 2022). This geometric perspective has led to efforts to estimate Local Intrinsic Dimensionality (LID) at the sample level (Stanczuk et al., 2024; Kamkari et al., 2024; Horvat & Pfister, 2024; Wenliang & Moran, 2023; Tempczyk et al., 2022), which has been used to characterize memorization (Ross et al., 2024; Ventura et al., 2024).
While our work builds on previous insights, it introduces key distinctions. Unlike approaches that define memorization at a global model level (Yoon et al., 2023; Gu et al., 2023), we focus on local, sample-wise behavior in probability densities. This aligns with the geometric perspective of (Ross et al., 2024; Bhattacharjee et al., 2023) but differs fundamentally by analyzing sharpness in the learned density rather than assuming an unobservable ground-truth distribution. We show that memorized samples consistently occupy high-sharpness regions throughout the denoising process, not just in the final density. Unlike LID-based methods (Ross et al., 2024), which apply only at the final step, our approach tracks memorization from the earliest stages of generation. This allows for early detection and intervention, offering a more proactive approach to mitigating memorization.
@@
Detecting and mitigating memorization during the generative process remains a challenging problem. Previous studies have explored various approaches to identify prompts that induce memorization in text-conditional DMs by comparing generated images to training data. For instance, Somepalli et al. (2023a) employed feature-based detectors like SSCD (Pizzi et al., 2022) and DINO (Caron et al., 2021), while Carlini et al. (2023) and Yoon et al. (2023) used calibrated $\ell_{2}$ distance in pixel space to quantify memorization, an approach later adopted by Stein et al. (2023). Additionally, Webster (2023) developed both white-box and black-box attacks, analyzing edges and noise patterns in generated images.
@@
Since memorization in DMs is often linked to specific text prompts, most mitigation strategies have focused on modifying prompts or adjusting attention mechanisms to reduce their influence (Wen et al., 2024; Ren et al., 2024; Ross et al., 2024). For example, Ross et al. (2024) rephrased prompts using GPT-4 to mitigate memorization. However, these interventions can degrade image quality or compromise user intent by altering model-internal components.
@@ 3 Preliminaries
In many applications the data $\mathbf{x}_{0}$ is often represented with an associated label $c$ (e.g., prompts or class labels). In these scenarios, the additional condition $c$ is incorporated into the model as $s_{\theta}(\mathbf{x}_{t},c)$, allowing it to estimate the score of the conditional density $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|c)\coloneqq s(\mathbf{x}_{t},c)$ via classifier free guidance Ho & Salimans (2021).
@@ 4.1 Memorization: Sharpness in Probability Landscape
While both sharpness and LID characterize curvature properties of probability density, LID is limited to analyzing sample space at $t\approx 0$, the final stage of generation. We extend memorization detection across all timesteps by leveraging sharpness via Hessian eigenvalues as a more versatile metric, enabling continuous monitoring throughout the generative process rather than relying solely on final output characteristics.
@@
We further validate our approach on Stable Diffusion Rombach et al. (2022), analyzing its $16,384$-dimensional latent space. Figure 2 reveals distinct patterns in both the number of non-negative eigenvalues and the magnitude of negative eigenvalues across different memorization categories (EM, PM, and non-memorized) at both initial and final sampling step. These patterns not only align with LID-based analysis at $t\approx 0$ but also demonstrate sharpness as a more generalizable memorization measure, capturing distinctive characteristics at generation onset.
@@ 4.2 Score Norm as a Sharpness Measure
A key observation is that the norm of the score function $\|s(\mathbf{x})\|$ inherently encodes information about the probability landscape’s curvature. In Gaussian distributions, the score norm is directly connected to the Hessian trace, as shown in the following result:
@@
While the distribution $\mathbf{x}_{t}$ in diffusion processes is not strictly Gaussian across all timesteps, recent findings indicate that for high to moderate noise levels (i.e., early to mid-stages in the reverse process), the learned score is dominated by the Gaussian component (Wang & Vastola, 2024). Moreover, in latent diffusion models where the data distributions are even more complex, the encoded latent variable $\mathbf{z}_{t}$ is explicitly regularized to stay near a Gaussian prior (Kingma, 2013; Rombach et al., 2022), reinforcing this approximation’s practical relevance.
@@ 4.4 Upscaling Eigenvalue Statistics via Hessian
This relationship, empirically verified in  Figure 4, demonstrates the Hessian-score product’s curvature-sensitive scaling effect. Building on this principle, we propose an enhanced version of Wen’s metric that improves early-stage sensitivity through second-order sharpness characterization:
@@ 4.5 Detecting Memorization in Stable Diffusion
For comparison, we implement six baseline methods. Carlini et al. (2023) analyzed generation density through pixel-wise $\ell_{2}$ distances across non-overlapping image tiles. Ren et al. (2024) detected memorized samples by identifying anomalous attention score patterns in text-conditioning during sampling. Chen et al. (2024) refined Wen et al. (2024)’s metric for partial memorization by incorporating end-token masks that empirically highlight locally memorized regions.
@@ Results.
Note that our metric is efficiently computed using Hessian-vector products without explicitly materializing the Hessian matrix, using automatic differentiation in libraries like PyTorch. As it requires only a single Hessian-vector product for detection, the computational overhead remains minimal.
@@ 5 Sharpness Aware Memorization Mitigation
Figure 6: Left: Comparing mitigation methods on SD v1.4 (top) and v2.0 (bottom) across five configurations. Lower SSCD indicates reduced memorization; higher CLIP shows better prompt alignment. Right: Qualitative comparison demonstrating SAIL’s effectiveness in preserving key image details (shown adjacent to the original image), whereas baseline methods exhibit quality degradation due to modified text conditioning. Images are generated using identical random seeds, with full prompts in Section D.2
@@ 5.1 Sharpness Aware Initialization Sampling
While $\|H_{\theta}^{\Delta}(\mathbf{x}_{T})s_{\theta}^{\Delta}(\mathbf{x}_{T})\|^{2}$ can be efficiently computed using Hessian-vector products, the gradient backpropagation required for optimization introduces computational overhead. To overcome the computational burden, we approximate the term using a Taylor expansion around $\mathbf{x}_{T}$:
@@
To evaluate mitigation strategies, we use the same memorized prompt set as in the detection experiments described in Section 4.5. However, since training images are required to assess the effectiveness of the mitigation methods, some prompts with inaccessible training images are excluded. Details are in Appendix D.
@@
For a fair comparison, all methods are evaluated using five distinct hyperparameter settings and optimized with the Adam optimizer at a learning rate of 0.05. For a detailed experimental settings, please refer to Section D.2.
@@ Results.
The advantage of SAIL is further illustrated in Figure 6 (right), where it generates images that preserve key prompt details, such as celebrity names or primary objects. In contrast, methods that modify text-conditional components often fail, as their approaches diminish the influence of such components during mitigation, potentially degrading user utility. Additional qualitative results of mitigation algorithms are provided in Appendix E.
@@ 6 Conclusion
We present a sharpness-based framework for detecting and mitigating memorization in diffusion models. By analyzing Hessian eigenvalues, we establish sharpness as a key indicator of memorization and introduce computationally efficient proxies using the Hessian trace and score norm. We further reinterpret Wen’s metric through this lens and propose Sharpness-Aware Initialization for Latent Diffusion (SAIL), an inference-time method that reduces memorization by selecting smoother initial noise. Experiments on 2D, MNIST, and Stable Diffusion confirm that our approach enables early detection and effective mitigation without degrading generation quality.
@@ A Gaussian log-density has
B.2 Proof of Lemma 4.2

### v3

# Understanding and Mitigating Memorization in Generative Models via Sharpness of Probability Landscapes

@@ 1 Introduction
Despite their successes, diffusion models suffer from memorization, where they replicate training samples instead of generating novel outputs (Carlini et al., 2023; Somepalli et al., 2023b; Webster, 2023). This issue is especially concerning when models are trained on sensitive data, leading to privacy risks (Orrick, 2023; Joseph Saveri, 2023). Addressing memorization is critical for ensuring the responsible deployment of generative models in real-world applications.
@@
Beyond detection, we introduce an inference-time mitigation strategy that reduces memorization by selecting initial diffusion noise from regions of lower sharpness. Our method, Sharpness-Aware Initialization for Latent Diffusion (SAIL), utilizes our sharpness metric to identify initializations that avoid trajectories leading to memorization. By simply adjusting the initial noise, SAIL steers the diffusion process toward smoother probability regions, mitigating memorization without requiring retraining. Unlike modifying the prompt, which can negatively affect generation quality, SAIL reduces memorization by carefully selecting the initial noise, enabling mitigation without changing the conditioning inputs.
@@
  • We provide a theoretical justification for the memorization detection metric introduced by Wen et al. (2024) through sharpness analysis.
  • We introduce a new sharpness measure that enables early-stage memorization detection during the diffusion process.
@@ 2 Related works
The memorization behavior of diffusion models (DMs) has been extensively studied (Somepalli et al., 2023b; Carlini et al., 2023; Wen et al., 2024), with prior work examining contributing factors such as prompt conditioning (Somepalli et al., 2023b), data duplication (Carlini et al., 2023; Somepalli et al., 2023a), and dataset size or complexity (Gu et al., 2023). Some studies have approached this issue from a geometric standpoint, drawing on the manifold learning conjecture (Fefferman et al., 2016; Pope et al., 2021), where exact memorization is associated with data points lying on a zero-dimensional manifold (Ross et al., 2024; Ventura et al., 2024; Pidstrigach, 2022).
This geometric perspective has led to efforts to estimate Local Intrinsic Dimensionality (LID) at the sample level (Stanczuk et al., 2024; Kamkari et al., 2024; Horvat & Pfister, 2024; Wenliang & Moran, 2023; Tempczyk et al., 2022), which has been used to characterize memorization (Ross et al., 2024; Ventura et al., 2024).
While our work is inspired by prior studies, it introduces several key distinctions. Unlike approaches that define memorization in terms of overall model behavior (Yoon et al., 2023; Gu et al., 2023), we focus on sample-specific behavior manifested in the learned probability density. Although our perspective is conceptually aligned with recent geometric interpretations (Ross et al., 2024; Bhattacharjee et al., 2023), our methodology diverges fundamentally by analyzing sharpness in the learned density, without relying on assumptions about an inaccessible ground-truth distribution. In contrast to manifold-based analyses that track variations in individual feature components (Ventura et al., 2024; Achilli et al., 2024), we show that sharpness, treated as an aggregated statistic, can be effectively estimated and used for detecting memorization. Moreover, unlike LID-based methods (Ross et al., 2024) that are restricted to the final denoising step, our approach reveals that memorized samples persistently occupy high-sharpness regions throughout the diffusion process. This allows for earlier detection and targeted intervention, enabling a more proactive and interpretable strategy for mitigating memorization.
@@
Detecting and mitigating memorization during the generative process remains a challenging problem. Previous studies have explored various approaches to identify prompts that induce memorization in text-conditional DMs by comparing generated images to training data. For instance, Somepalli et al. (2023a) employed feature-based detectors like SSCD (Pizzi et al., 2022) and DINO (Caron et al., 2021), while Carlini et al. (2023) and Yoon et al. (2023) used calibrated $\ell_{2}$ distance in pixel space to quantify memorization. Webster (2023) developed both white-box and black-box attacks, analyzing edges and noise patterns in generated images.
@@
Since memorization in DMs is often linked to specific text prompts, most mitigation strategies have focused on modifying prompts or adjusting attention mechanisms to reduce their influence (Wen et al., 2024; Ren et al., 2024; Ross et al., 2024). For example, Ross et al. (2024) rephrased prompts using GPT-4 to mitigate memorization. However, these interventions often degrade image quality or compromise user intent by altering model-internal components.
@@ 3 Preliminaries
In many applications the data $\mathbf{x}_{0}$ is often represented with an associated label $c$ (e.g., prompts or class labels). In these scenarios, the additional condition $c$ is incorporated into the model as $s_{\theta}(\mathbf{x}_{t},c)$, allowing it to estimate the score of the conditional density $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|c)\coloneqq s(\mathbf{x}_{t},c)$ via classifier free guidance (Ho & Salimans, 2021).
@@ 4.1 Memorization: Sharpness in Probability Landscape
While both sharpness and LID characterize curvature properties of probability density, LID is limited to analyzing sample space at $t\approx 0$, where the generated image emerges. In contrast, we extend memorization detection across all timesteps by leveraging sharpness via Hessian eigenvalues as a more versatile metric, enabling continuous monitoring throughout the generative process rather than relying solely on final output characteristics.
@@
We further validate our approach on Stable Diffusion (Rombach et al., 2022), analyzing its $16,384$-dimensional latent space. Figure 2 reveals distinct patterns in both the number of non-negative eigenvalues and the magnitude of negative eigenvalues across different memorization categories (EM, PM, and non-memorized) at both initial and final sampling step. These patterns not only align with LID-based analysis at $t\approx 0$ but also demonstrate sharpness as a more generalizable memorization measure, capturing distinctive characteristics at generation onset.
@@ 4.2 Score Norm as a Sharpness Measure
A key observation is that the norm of the score function $\|s(\mathbf{x})\|$ inherently encodes information about the probability landscape’s curvature. In Gaussian distributions, the score norm is directly connected to the Hessian trace, as shown in the following result. (Appendix State).
@@
This result extends to non-Gaussian distributions under mild regularity assumptions (Appendix State). For theoretical clarity and ease of analysis, however, we focus on the Gaussian case. While the distribution $\mathbf{x}_{t}$ in diffusion processes is not strictly Gaussian at every timestep, recent studies show that at moderate to high noise levels, corresponding to the early and middle stages of the reverse process—the learned score is predominantly governed by its Gaussian component (Wang & Vastola, 2024). This approximation is further justified in latent diffusion models, where the latent variable $\mathbf{z}_{t}$ is explicitly regularized toward a Gaussian prior (Kingma, 2013; Rombach et al., 2022), despite the complexity of the original data distribution.
@@ 4.4 Upscaling Eigenvalue Statistics via Hessian
This relationship, empirically verified in  Figure 4, demonstrates the curvature-sensitive scaling effect of the Hessian score product. Building on this principle, we propose an enhanced version of Wen’s metric that improves early-stage sensitivity through second-order sharpness characterization:
@@ 4.5 Detecting Memorization in Stable Diffusion
For comparison, we implement six baseline methods. Among them, Carlini et al. (2023) analyzed generation density by measuring pixel-wise $\ell_{2}$ distances across non-overlapping image tiles, aiming to detect memorized content based on local similarity patterns. Ren et al. (2024) detected memorized samples by identifying anomalous attention score patterns in text-conditioning during sampling. Chen et al. (2024) refined Wen et al. (2024)’s metric for partial memorization by incorporating end-token masks that empirically highlight locally memorized regions.
@@ Results.
Importantly, our metric can be efficiently computed using Hessian-vector products without explicitly forming the full Hessian matrix. Leveraging automatic differentiation frameworks such as PyTorch, a single Hessian-vector product suffices for detection, incurring minimal overhead.
@@ 5 Sharpness Aware Memorization Mitigation
Figure 6: Left: Comparison of inference-time mitigation methods on SD v1.4 (top) and v2.0 (bottom), evaluated across five hyperparameter configurations per method. Lower SSCD scores indicate reduced memorization, while higher CLIP scores show better prompt-image alignment. Right: Qualitative comparison demonstrating SAIL’s effectiveness in preserving key image details (shown adjacent to the original image), whereas baseline methods exhibit quality degradation due to modified text conditioning. Images are generated using identical random seeds, with full prompts in Section D.2
@@ 5.1 Sharpness Aware Initialization Sampling
While $\|H_{\theta}^{\Delta}(\mathbf{x}_{T})s_{\theta}^{\Delta}(\mathbf{x}_{T})\|^{2}$ can be efficiently computed using Hessian-vector products, the gradient backpropagation required for optimization introduces computational overhead. To overcome the burden, we approximate the term using a Taylor expansion around $\mathbf{x}_{T}$:
@@
To evaluate mitigation strategies, we use the same memorized prompt set employed in the detection experiments described in Section 4.5. However, since verifying mitigation effects requires access to training images, we exclude prompts whose corresponding training samples are unavailable. Further details are in Appendix D.
@@
For a fair comparison, all methods are evaluated using five distinct hyperparameter settings and optimized with the Adam optimizer at a learning rate of 0.05. For a detailed experimental settings, refer to Section D.2.
@@ Results.
The advantage of SAIL is evident in Figure 6 (right), where it generates images that faithfully preserve key prompt details, such as celebrity names and primary objects. In contrast, methods that modify text-conditional components often reduce the influence of those components during mitigation, leading to degraded alignment with the original prompt and potentially diminishing user utility. Additional qualitative results for algorithms are provided in Appendix E.
@@ 6 Conclusion
We propose a sharpness-based framework for detecting and mitigating memorization in diffusion models. Our analysis identifies Hessian-based sharpness as a reliable indicator of memorization and introduces an efficient proxy based on the score norm. This perspective also provides a theoretical interpretation of the memorization detection metric proposed by Wen et al. (2024). Building on this foundation, we introduce Sharpness-Aware Initialization for Latent Diffusion (SAIL), an inference-time method that reduces memorization by selecting low-sharpness initial noise. Experiments on synthetic 2D data, MNIST, and Stable Diffusion demonstrate that our approach enables early detection and effective mitigation, all without degrading generation quality.
Acknowledgement
This work was supported in part by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2024-00457882, AI Research Hub Project), the Ministry of Science and ICT (MSIT), South Korea, under the Information Technology Research Center (ITRC) Support Program (IITP-2025-RS-2022-00156295), and IITP grant funded by the Korean Government (MSIT) (No. RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)).
Impact Statement
Our work aims to advance the understanding and mitigation of memorization in diffusion models, a phenomenon closely tied to potential privacy risks. By proposing a framework to detect and reduce memorization, we seek to enhance the responsible deployment of generative models, especially when they are trained on sensitive data. This approach could contribute positively by lowering the risk of unintentionally revealing private information.
@@ A Gaussian log-density has
This result generalizes to non-Gaussian distributions under weak regularity conditions (Hyvärinen, 2005). Although we chose the Gaussian assumption to facilitate theoretical extensions and applications, we will still present the original generalization here.
B.2 Generalization of Lemma 4.1
State.
For a random vector $\mathbf{x}\sim p(\mathbf{x})$ with regularity conditions $\mathbb{E}\left[||s(\mathbf{x})||^{2}\right]<\infty$ and $\underset{\|x\|\to\infty}{\lim}p(\mathbf{x})s(\mathbf{x})=\mathbf{0}$,
${\mathbb{E}\left[\|s(\mathbf{x})\|^{2}\right]}=-\mathbb{E}\left[\textup{{{tr}}}(H(\mathbf{x}))\right].$
Proof.
Write $s_{i}(\mathbf{x})=\partial_{x_{i}}\log p(\mathbf{x})$. Because $s_{i}\,p=\partial_{x_{i}}p$,
$\mathbb{E}\!\bigl{[}\|s(\mathbf{x})\|^{2}\bigr{]}=\sum_{i=1}^{d}\!\int s_{i}(\mathbf{x})\,\partial_{x_{i}}p(\mathbf{x})\,d\mathbf{x}.$
For each $i$ integrate by parts:
$\int s_{i}\,\partial_{x_{i}}p=\int\partial_{x_{i}}\!\bigl{[}p\,s_{i}\bigr{]}\,d\mathbf{x}-\int p\,\partial_{x_{i}}s_{i}\,d\mathbf{x}.$
The first term is a surface integral over the sphere of radius $R$; by the assumed boundary condition it vanishes as $R\to\infty$. Hence $\int s_{i}\,\partial_{x_{i}}p=-\int p\,\partial_{x_{i}}s_{i}$. Summing over $i$ gives
$\mathbb{E}\!\bigl{[}\|s(\mathbf{x})\|^{2}\bigr{]}=-\int p(\mathbf{x})\,\sum_{i=1}^{d}\partial_{x_{i}}s_{i}(\mathbf{x})\,d\mathbf{x}=-\mathbb{E}\!\bigl{[}\textup{{{tr}}}(H(\mathbf{x}))\bigr{]}.$
B.3 Proof of Lemma 4.2
@@ D.2 Memorization Mitigation
As the metric proposed by Wen et al. (2024) also captures sharpness, one may consider replacing the first term of $\mathcal{L}_{\text{SAIL}}(\mathbf{x}_{T})$ with $\lVert s_{\theta}(\mathbf{x}_{t},c)-s_{\theta}(\mathbf{x}_{t})\rVert^{2}.$ However, we empirically find that this alternative fails to converge and is therefore ineffective for mitigation. This may be due to the higher sensitivity of our proposed metric during the initial phase of generation.